OpenStudy (anonymous):

Solve for x. 9(x + 1) = 25 + x

OpenStudy (anonymous):

x = 2 x = 3 x = 4 x = 5

OpenStudy (anonymous):

x=3

OpenStudy (igreen):

Distribute 9..do you know how to do that? @Olayshia22

OpenStudy (igreen):

@cole1117 Please do not give answers, it's against OpenStudy's Code-of-Conduct. We are here to guide the asker to the answer, so we need his/her participation. Please read the \(\bf\color{blue}{\underline{{\href{http:///openstudy.com/code-of-conduct}{Code-of-Conduct}}}}\) to learn more.

OpenStudy (swamijatin2):

9(x+1)=25+x 9x+9=25+x 9x-x=25-9 8x=16 x=16/8 x=2

OpenStudy (igreen):

@swamijatin2 Please do not give answers, it's against OpenStudy's Code-of-Conduct. We are here to guide the asker to the answer, so we need his/her participation. Please read the \(\bf\color{blue}{\underline{{\href{http:///openstudy.com/code-of-conduct}{Code-of-Conduct}}}}\) to learn more.

OpenStudy (swamijatin2):

sorry for that

OpenStudy (anonymous):

My bad, I was just about to start with the 9x + 1= 25 + x then let him go from there

OpenStudy (igreen):

That would be 9x + 9 = 25 + x

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

hit the wrong number. It's been a long morning.

OpenStudy (anonymous):

Thank you